# Quantification of Interdependent Emotion Dynamics in Online Interactions

### v2

@@ Discussion
We find that live videos, while central to content delivery, have limited direct influence on users’ emotional expressions. Among 92,412 live chat messages on YouTube, emotional dynamics are primarily shaped by user interactions rather than passive video consumption, with peer-driven effects outweighing video influence by a factor of four. This supports the view that emotions are inherently communicative, emerging through social interaction rather than passive reception [72].
@@
Beyond general social influence, emotional mimicry and social appraisal processes further explain the primacy of peer interactions. Emotional mimicry depends on contextual interpretation rather than direct replication of expressions [77]. On YouTube Live, video content provides the contextual backdrop, but emotions propagate mainly through user interactions. Similarly, emotional expressions function as social signals, helping individuals navigate ambiguity by inferring meaning from others’ emotions [78]. In this setting, chat-based emotional expressions guide audience engagement with video content. This aligns with Affect Theory of Social Exchange [79], which holds that emotions emerging from social interactions reinforce group cohesion and collective experiences.
Together, these findings suggest that YouTube Live fosters a socially constructed emotional environment, where peer interactions - not the video itself - drive emotional expression. Consequently, as language models grow more sophisticated, the threat of bots mimicking human users to endogenously manipulate collective emotions becomes an increasing concern [80].
@@
First, anger and disgust strongly reinforce one another. Their reciprocal triggering, along with high self-excitation, aligns with the role of these moral emotions in driving moral outrage. Previous work has shown that co-occurrence of anger and disgust is predictive of moral outrage [88]; our findings suggest this co-occurrence may be further amplified by mutual excitation. This mechanism could help explain the frequent escalation of outrage in online discourse and its potential contribution to polarization and social fragmentation [89, 90]. We note, however, that part of this effect may be due to the frequent co-labeling of anger and disgust in the same message (see SI Appendix).
@@
Limitations
A key limitation of our analysis is that we approximate the emotional content of the video solely through its subtitles. While this text-based approach captures the semantic cues of emotional expression, it omits prosodic and visual signals such as tone of voice, facial expressions, and body language. Previous studies have shown that discrete emotion measures derived from text correlate with self-reported emotions, but may diverge from facial and vocal indicators [48]. Future work integrating multimodal data—including audio and visual cues—could improve emotion inference and offer a more complete picture of how emotional dynamics unfold across communication channels.
@@
Finally, our modeling framework is based on a discrete event Hawkes process that treats emotions as categorical states. While this setup easily generalizes to more than six discrete emotions, it does not directly support alternative representations that view emotions as points in a continuous valence–arousal space [96]. To capture such continuous affective states, the framework could be extended to a marked Hawkes process, where each event carries additional information or marks - such as a real-valued vector indicating valence and arousal. This extension would allow the model to capture not only the timing of events but also their qualitative characteristics, making it well suited to reflect the graded nature of emotional responses in continuous spaces.
@@ Conclusions
We analyzed 92,412 YouTube Live chat messages from 397 videos to study emotional contagion in social networks and livestream platforms. Our findings show that user interactions are four times more emotionally influential than passive video content. Joy is three times more contagious than negative emotions, but negative emotions persist nearly twice as long, potentially creating lasting negative moods. We also observed cross-excitation effects, where negative posts can trigger positive emotions in trolls, revealing antisocial dynamics that harm community cohesion. These insights have practical implications: promoting joyful content and enabling active communication (e.g., live chats or bots) can foster emotional engagement and positivity. However, the lingering impact of negative emotions highlights the need for timely moderation. Platforms must balance emotional amplification with community well-being. Designing for positive contagion while managing toxic interactions is key to sustaining healthy online environments. Future research could explore cultural influences and long-term impacts of emotional contagion strategies.
@@ Appendix A Summary Statistics of YouTube Live Chat Data
We show some additional summary statistics of our dataset. Once we filtered the data by removing inactive videos, we compiled 92,412 live chat messages from 397 videos, amassing a total of 780 hours of material and encompassing 27 distinct topics. The distribution of topics is shown in Figure A.1. The distribution of emotional events is shown in Figure A.2. We also show the distribution of the median time between each event arrival in Figure A.3. Since we label emotions non-exclusively, we show the co-occurrence patterns across 6 basic emotions in Figure A.4. In general, negative emotions have a higher tendency to co-occur with one another. In particular, anger and disgust have a relatively high co-occurrence.
@@
Appendix B Robustness Analysis on Video Influence
Figure B.1: To test the validity of our video parametrization method, we repeat the analysis while changing the parameters of the video influence function as well as the influence function shape. This Figure shows the estimated parameters of our log-likelihood function in Equation (5) in the main text. Figures 1(a) and 1(b) show the estimated parameters when multiplying the maximum and median in the log-normal function by a factor of 0.5 and 2, respectively. Figure 1(c) shows the estimated parameters when parametrizing the video influence with a power law shape. The estimated parameters are highly consistent with our main result (Figure 2 in the main text).
Figure B.2: Figures 1(a) and 1(b) show the relative source of influence when multiplying the maximum and median in the log-normal function by a factor of 0.5 and 2, respectively. Figure 1(c) shows the relative source of influence when parametrizing the video influence with a power law shape. The left column shows the relative percentage of endogenous and exogenous influence. The right column zooms in on the exogenous influence and examines the relative source of mother events from spontaneous user expressions and video influence. The relative influences are highly consistent with our main result (Figure 3 in the main text). The endogenous influence is comparably larger than the exogenous influence, and the source of mother events is predominately spontaneous user expressions.
To examine the impact of video content on live chat discussions, we have parameterized and interpolated the prevalence of emotions in the video via sums of log-normal shapes (see Methods). Here, we test the robustness of our results with respect to changes in the log-normal parameters. We have defined the shape of the log-normal function by calibrating its mode (maximum value), and the median. We assumed that the intensity of each transcript peaks $2$ seconds after appearance such that the maximum of the log-normal function lies at $2$ seconds. We further assumed that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance such that the median is equal to $10$ seconds. Here we repeat our analysis while changing the log-normal function maximum and median by factors of 0.5 and 2 to test the robustness of our results. This effectively changes assumptions about the audience’s reaction time and memory of prior content. We observe in Figures 1(a), 1(b), 2(a), and 2(b) that the results are qualitatively consistent with the results in the main text. This suggests that our results aren’t particularly sensitive to the parametrization of the video content.
Human memory decays following an approximate power law [67]. The log-normal function resembles the shape of a power law over a limited range of the variable, which is all the larger, the larger is $\sigma$ [68]. We thus test our results with an alternative parametrization of the video influence in a power law shape where $s^{f}_{\tau_{j}}(t)=\frac{1}{(t-\tau_{j})^{c}}$. We show examples in Figure 1(c) and 2(c) for $c=2.5$, and find that the results are qualitatively consistent with the results we obtain from the log-normal over a range of $c$ values from $2.5$ to $10$.
Appendix C Dynamics of 4 Emotions
Figure C.1: This figure shows the estimated parameters of our log-likelihood function in Equation (5) in the main text for 4 emotions. The estimations are consistent with our main results (Figure 2 in the main text).
Figure C.2: This left plot shows the relative ratios of endogenous and exogenous influences. The right plot shows the ratios of mother events from spontaneous expression and video influence. We observe that the endogenous influence dominates the emotion dynamics in YouTube live chats. In addition, the mother events mostly stem from spontaneous user expressions rather than video influence. This is consistent with our main result (Figure 3 in the main text).
Our main analysis models the commonly studied 6 basic emotions [50]. Here we test the robustness of our results by modeling $|\mathcal{E}|=4$ emotions: joy, anger, disgust, and sadness, removing fear and surprise because they have small excitatory patterns and low overall intensity. We obtain a sample of 288,581 live chat messages across 1,405 videos. Reducing the dimensionality of the multivariate Hawkes process reduces the number of parameters to be estimated from 84 to 40. We obtain our results for 4 emotions in Figures C.1 and C.2. We find that the dynamics of 4 emotions are qualitatively similar to our main results on 6 emotions (Figures 2 and 3 in the main text).
Appendix D Sub-sample Analysis with Different Video Types
Figure D.1: This figure shows the estimated parameters of our log-likelihood function in Equation (5) in the main text for sports videos. The estimations are largely consistent with our main result (Figure 2 in the main text).
Figure D.2: This left plot shows the relative ratios of endogenous and exogenous influences for sports videos. The right plot shows the ratios of mother events from spontaneous expression and video influence for sports videos. We observe that the exogenous influence has a larger impact on emotion intensities for sports videos compared to our main result (Figure 3 in the main text). In particular, the emotions of surprise and fear are largely exogenously driven. The mother events are induced more by spontaneous expressions rather than video influence.
Figure D.3: This figure shows the estimated parameters of our log-likelihood function in Equation (5) in the main text for political videos. We observe that joy remains the most contagious emotion, and the cross-triggering effects between anger and disgust are prominent in the political context. The emotions of surprise and fear impose significantly longer direct memories on users.
Figure D.4: This left plot shows the relative ratios of endogenous and exogenous influences for political videos. The right plot shows the ratios of mother events from spontaneous expression and video influence for political videos. We observe that endogenous influence strongly dominates the emotion dynamics in political discussions. As opposed to our main result in (Figure 3 in the main text), political videos are more stimulative of discussions than the average video content.
Our main analysis examines general emotional and behavioral patterns among online users by compiling 92,412 YouTube live chat messages across 27 topics. Yet the susceptibility of emotion dynamics to video influence and peer interactions is potentially subject to the type of video and the audience community. Here, we conduct sub-sample analyses on representative video types to identify heterogeneity in the evolution of online emotion dynamics. First, we focus on live sports videos with a total of 11,172 live chat messages. Figures D.1 and D.2 show estimated parameters and relative sources of influence for live sports shows. We identify similarities to our results in the main text. The emotion of joy is the most contagious. Anger tends to self-excite and trigger disgust. Given the natural conflict from sports competitions, we observe cross-triggering between sadness and joy. In particular, users have a much longer direct memory of fear than other emotion types. Users have a much higher propensity for joyful expression which coincides with the entertaining nature of sports events. However, compared to our main result (Figure 3 in the main text), the emotion dynamics are more susceptible to exogenous influences. In particular, the emotions of surprise and fear are the most exogenously driven.
Second, we study political videos with a total of 3,043 live chat messages to understand how emotions evolve in political discussions. We highlight the cross-triggering of anger and disgust which contributes to moral outrage commonly observed in the political context. Users have longer-lasting direct memories of fear and surprise in online discussions. Political content is often accompanied by contentious discussions. In fact, we observe that endogenous interactions dominate exogenous influence by a large margin across emotion types. However, different from our main result (Figure 3 in the main text), the video content has a comparatively larger influence than spontaneous expressions in driving the emotion dynamics. This shows that political videos are highly stimulative in evoking emotional reactions in the audience. We observe that heterogeneity in online emotion evolution can arise from differences in video content and audience characteristics.
Appendix E Dynamics of Live Chats with No Emotional Content
Parameters: $\alpha$ $\gamma$ $\mu$ $\nu$
Mean 0.857 1.967 0.150 0.029
Standard Deviation 0.018 0.167 0.022 0.006
Appendix F Baseline Intensity Calibration in Hawkes Process
@@
Appendix G Log-Likelihood Derivation
@@
Appendix H Testing Log-Likelihood Fits with Synthetic Data
@@
We estimate parameter values that maximize the log-likelihood function (G) via Quasi-Newton optimization. Without loss of generality, we normalize the log-likelihood function by the number of events $N^{e}$, in order to compare the log-likelihood values across different sample sizes. We use the Python ticks library 222 https://x-datainitiative.github.io/tick/index.html to simulate Hawkes processes with parameters that we refer to as the ground-truth. Subsequently, we compare the estimated parameters with their ground-truth values to evaluate performance. Since the parameters for each emotion are estimated separately, without loss of generality, we only show the synthetic test results for one emotion (process). The other processes have comparable performances. To assess the robustness of the estimation, we bootstrap the data and repeat the estimation 100 times.
@@
Figure H.1: We plot the shape of the log-likelihood function in Equation (G) varying the target parameter while holding the other parameters fixed at their ground-truth values. The ground-truth values of the target parameter are indicated by the red dots. The log-likelihood values are normalized by the number of events $N^{e}$ to be comparable across varying numbers of data points. The inset plots show the distribution of the fitted parameters from 100 bootstrapped samples. The red dashed lines indicate the ground-truth values for the parameter of interest. We show results for the first process in a 6-dimensional multivariate Hawkes process. The other processes have comparable performances.

### v3

@@ Discussion
We find that live videos, while central to content delivery, have comparably less direct influence on users’ emotional expressions than peer interactions.
Among 92,412 live chat messages on YouTube, emotional dynamics are disproportionately shaped by user interactions rather than passive video consumption, with peer-driven effects outweighing video influence by a factor of four. This supports the view that emotions are inherently communicative, emerging primarily through social interaction rather than passive reception [72].
@@
Beyond general social influence, emotional mimicry and social appraisal processes further explain the primacy of peer interactions. Emotional mimicry depends on contextual interpretation rather than direct replication of expressions [77]. On YouTube Live, video content provides the contextual backdrop, but emotions propagate comparably more through user interactions. Similarly, emotional expressions function as social signals, helping individuals navigate ambiguity by inferring meaning from others’ emotions [78]. In this setting, chat-based emotional expressions guide audience engagement with video content. This aligns with Affect Theory of Social Exchange [79], which holds that emotions emerging from social interactions reinforce group cohesion and collective experiences.
Together, these findings suggest that YouTube Live fosters a socially constructed emotional environment, where peer interactions disproportionately drive emotional expression compared to the video content. Consequently, as language models grow more sophisticated, the threat of bots mimicking human users to endogenously manipulate collective emotions becomes an increasing concern [80].
@@
First, anger and disgust strongly reinforce one another. Their reciprocal triggering, along with high self-excitation, aligns with the role of these moral emotions in driving moral outrage. Previous work has shown that co-occurrence of anger and disgust is predictive of moral outrage [88]; our findings suggest this co-occurrence may be further amplified by mutual excitation. This mechanism could help explain the frequent escalation of outrage in online discourse and its potential contribution to polarization and social fragmentation [89, 90]. We note, however, that part of this effect may be due to the frequent co-labeling of anger and disgust in the same message (see Appendix A).
@@
Limitations and Future Work
A key limitation of our analysis is that we approximate the emotional content of the video solely through its subtitles. While this text-based approach captures the semantic cues of emotional expression, it omits prosodic and visual signals such as tone of voice, facial expressions, and body language. Previous studies have shown that discrete emotion measures derived from text correlate with self-reported emotions, but may diverge from facial and vocal indicators [48]. Future work integrating multimodal data, including audio and visual cues, could improve emotion inference and offer a more complete picture of how emotional dynamics unfold across communication channels.
@@
Another limitation of the present study is that emotions are modeled as discrete categories of equal magnitude. Although based on the widely used six basic emotions model [50], it does not account for variations in emotional intensity or the continuous structure of affective experience. An immediate extension would be to adopt a marked Hawkes process, in which each emotional event is associated with a continuous-valued mark reflecting its strength or salience [96]. This could be achieved by learning to assign each emotional expression a continuous intensity score, or, presumably even more insightfully, by adopting a continuous representation of emotions grounded in the circumplex model of affect. In this model, emotions are embedded in a two- or three-dimensional space defined by axes such as valence (positive vs. negative affect), arousal (high vs. low activation), and dominance (control vs. submission) [97, 98]. A marked Hawkes process defined over such a space would enable the modeling of how emotional expressions not only evolve in frequency but also drift through affective space over time. This would allow us to capture more fine-grained patterns of emotional contagion, such as whether high arousal is more contagious than low arousal, or whether emotional dynamics differ along the dominance axis.
In parallel, future work could relax the assumption of additivity inherent in classical Hawkes models. One particularly promising extension is the Quadratic Hawkes process which incorporates both linear and quadratic feedback terms [59, 99]. This allows the intensity of future emotional expressions to depend not only on the sum of past events, but also on pairwise combinations of past emotions. Such a framework naturally accommodates non-linear interactions, e.g., whether the co-occurrence of surprise and sadness increases the likelihood of subsequent anger, and even allows for inhibition, as negative-valued quadratic terms can suppress future intensities. By combining the continuous circumplex representation of emotions with a marked Quadratic Hawkes process, future research could investigate how different dimensions of affect interact to shape emotional dynamics at the group level in real time.
Together, these extensions offer a powerful and flexible modeling framework to capture the rich, non-linear, and multi-dimensional nature of emotion contagion in online and offline interactions.
@@ Conclusions
We analyzed 92,412 YouTube Live chat messages from 397 videos to study emotional contagion in social networks and livestream platforms. Our findings show that user interactions are four times more emotionally influential than passive video content. Joy is three times more contagious than negative emotions, but negative emotions persist nearly twice as long, potentially creating lasting negative moods. We also observed cross-excitation effects, where negative posts can trigger positive emotions in trolls, revealing antisocial dynamics that harm community cohesion. These insights have practical implications: promoting joyful content and enabling active communication (e.g., live chats or bots) can foster emotional engagement and positivity. However, the lingering impact of negative emotions highlights the need for timely moderation. Platforms must balance emotional amplification with community well-being. Designing for positive contagion while managing toxic interactions is key to sustaining healthy online environments. Future research could explore cultural influences and the long-term impacts of emotional contagion strategies.
Data Availability Statement
Publicly available data are obtained from YouTube. Data needed to replicate this study are available at https://github.com/ivylyis/Quantification-of-the-Self-Excited-Emotion-Dynamics-in-Online-Interactions.
Code Availability Statement
Code needed to replicate this study, in Python 3.7.6, is available at https://github.com/ivylyis/Quantification-of-the-Self-Excited-Emotion-Dynamics-in-Online-Interactions.
@@ Appendix A Summary Statistics of YouTube Live Chat Data
We show some additional summary statistics of our dataset. Once we filtered the data by removing inactive videos, we compiled 92,412 live chat messages from 397 videos, amassing a total of 780 hours of material and encompassing 27 distinct topics. The distribution of topics is shown in Figure A.1. The distribution of emotional events is shown in Figure A.2. We also show the distribution of the median time between each event arrival in Figure A.3. To exemplify our emotion labeling approach, we showcase representative live chat messages and predicted emotions in Table A.1. Since we label emotions non-exclusively, we show the co-occurrence patterns across 6 basic emotions in Figure A.4. In general, negative emotions have a higher tendency to co-occur with one another. In particular, anger and disgust have a relatively high co-occurrence.
To examine whether expressing different emotions requires systematically different levels of effort, we analyze the distribution of message lengths across emotion categories. As shown in Figure A.5, the variation in text length within each category is substantially larger than the differences between categories. This suggests that observed asymmetries in emotional dynamics between positive and negative emotions are unlikely to be driven by differences in text length or typing effort.
@@
Figure A.5: Distribution of YouTube Live Chat message lengths across emotion categories. The whiskers indicate the 5th to 95th percentiles, with both the median and mean marked within each box. The variation in text length within each emotion category is substantially larger than the differences in average text length across categories, suggesting that message length is not systematically biased by emotion type.
Table A.1: Sample live chat messages and detected emotions
Live-chat comment Emotion(s)
Thanks so much for doing this. I feel like I got to go on an adventure during my lunch break! joy
That is unbelievable surprise
I hate laugh tracks!! Some of the shows on Netflix use it and it drives me nuts! anger, disgust
I can’t stand haters on high horses reporting people that are not even doing anything morally wrong anger, disgust
His voice sounds scary fear
So sorry, don’t cry anymore sadness
What time is it no emotions
Appendix B Baseline Intensity Calibration in Hawkes Process
@@
Appendix C Log-Likelihood Derivation
@@
Appendix D Testing Log-Likelihood Fits with Synthetic Data
@@
We estimate parameter values that maximize the log-likelihood function (C) via Quasi-Newton optimization. Without loss of generality, we normalize the log-likelihood function by the number of events $N^{e}$, in order to compare the log-likelihood values across different sample sizes. We use the Python ticks library 333 https://x-datainitiative.github.io/tick/index.html to simulate Hawkes processes with parameters that we refer to as the ground-truth. Subsequently, we compare the estimated parameters with their ground-truth values to evaluate performance. Since the parameters for each emotion are estimated separately, without loss of generality, we only show the synthetic test results for one emotion (process). The other processes have comparable performances. To assess the robustness of the estimation, we bootstrap the data and repeat the estimation 100 times.
@@
Figure D.1: We plot the shape of the log-likelihood function in Equation (C) varying the target parameter while holding the other parameters fixed at their ground-truth values. The ground-truth values of the target parameter are indicated by the red dots. The log-likelihood values are normalized by the number of events $N^{e}$ to be comparable across varying numbers of data points. The inset plots show the distribution of the fitted parameters from 100 bootstrapped samples. The red dashed lines indicate the ground-truth values for the parameter of interest. We show results for the first process in a 6-dimensional multivariate Hawkes process. The other processes have comparable performances.